# How segmented is my network?

### v2

@@ I Introduction
Network segmentation is a widely recommended security practice for limiting lateral movement and reducing the impact of breaches. Modern security frameworks routinely advise practitioners to divide networks into smaller, isolated zones and tightly control communication between them. Despite this emphasis, security teams lack a simple way to assess how segmented a network is. In practice, segmentation is usually assessed qualitatively using architecture diagrams, policy reviews, or expert judgment.
The lack of a quantitative measure has concrete consequences. Organizations cannot objectively compare segmentation across business units, validate that architecture changes improved isolation, or determine whether network complexity is justified by security benefits. To our knowledge, no prior work has provided a single, interpretable scalar metric for segmentedness—leaving practitioners unable to objectively measure what security frameworks universally recommend. Further, practitioners are left without objective evidence to support segmentation decisions or investments.
@@ III-E Expected Number of Reachable Neighbors
$|N(v)|=\sum_{u\in V\setminus\{v\}}X_{u}.$ (4)
For a fixed node $v$ and a uniformly random choice of another node $u\neq v$, the probability that $\{u,v\}\in E$ equals $F(G)$, since $F(G)$ represents the fraction of all node pairs for which communication is permitted. Therefore,
$\mathbb{E}[X_{u}]=F(G)\quad\text{for all }u\neq v.$ (5)
By linearity of expectation,
$\mathbb{E}[|N(v)|]=\sum_{u\in V\setminus\{v\}}\mathbb{E}[X_{u}]=(n-1)\,F(G).$ (6)
@@ VI-B Estimation over Networks with Communities
In the SBM with $K$ equal-sized blocks, a graph is generated on $n$ labeled nodes by first partitioning the nodes into $K$ disjoint groups of equal size. Edges between nodes in the same block are included independently with probability $p_{\text{in}}$, while edges between nodes in different blocks are included independently with probability $p_{\text{out}}$. Although all edge decisions remain independent, the probability of an edge now depends on the block memberships of the node pair. The resulting global edge density is a weighted average of $p_{\text{in}}$ and $p_{\text{out}}$:
@@ Appendix A Graph-based Measures’ Comparison
Table V shows all three graph-based measures from Section II-C for the ten densest networks in the Cisco Secure Workload dataset. The absolute correlation coefficient calculated over this table between Flatness and the Fiedler Value is $0.136$; that between Flatness and Modularity is $0.568$. The low correlation with Fiedler Value ($0.136$) and thr moderate correlation with Modularity ($0.568$) suggests the three measures capture related but distinct aspects of networks (graphs).
@@ Appendix B Bayesian Inference for High-segmentedness
To address this, we introduce a Bayesian Beta-binomial model to provide a more robust estimation of uncertainty at the boundaries [11]. The Beta distribution is the conjugate prior for the binomial process, making it an ideal choice for updating our beliefs about network flatness based on sampled data.

### v3

@@ I Introduction
Network segmentation is a widely recommended security practice for limiting lateral movement and reducing the impact of breaches. Modern security frameworks routinely advise practitioners to divide networks into smaller, isolated zones and tightly control communication between them. Despite this emphasis, security teams lack a simple way to assess how segmented a network is. In practice, segmentation is usually assessed qualitatively using architecture diagrams, network security policy reviews, or expert judgment. This gap exemplifies a broader challenge identified by the 2025 National Academies Cyber Hard Problems report: the absence of security metrics that are reliable, quantifiable, and repeatable [20].
The lack of a quantitative measure has concrete consequences. Organizations cannot objectively compare segmentation across business units, validate that architecture changes improved isolation, or determine whether network complexity is justified by security benefits. To our knowledge, no prior work has provided a single, interpretable scalar metric for segmentedness—leaving practitioners unable to objectively measure a practice that security frameworks universally recommend. Further, practitioners are left without objective evidence to support segmentation decisions or investments.
@@ III-E Expected Number of Reachable Neighbors
$|N(v)|=\sum_{u\in V\setminus\{v\}}X_{u}=\deg(v).$ (4)
Since $v$ is selected uniformly at random from $V$, the expected value of $|N(v)|$ equals the average node degree in $G$:
$\mathbb{E}[|N(v)|]=\frac{1}{n}\sum_{w\in V}\deg(w)=\frac{2|E|}{n}.$ (5)
Substituting $|E|=\binom{n}{2}F(G)=\frac{n(n-1)}{2}F(G)$ yields
$\mathbb{E}[|N(v)|]=(n-1)\,F(G).$ (6)
@@ VI-B Estimation over Networks with Communities
In the SBM with $K$ equal-sized blocks, a graph is generated on $n$ labeled nodes by first partitioning the nodes into $K$ disjoint groups of equal size. Edges between nodes in the same block are included independently with probability $p_{\text{in}}$, while edges between nodes in different blocks are included independently with probability $p_{\text{out}}$. Although all edge decisions remain independent, the probability of an edge now depends on the block memberships of the node pair. The resulting global edge density is a weighted average of $p_{\text{in}}$ and $p_{\text{out}}$ for large $n$:
@@ Appendix A Graph-based Measures’ Comparison
Table V shows all three graph-based measures from Section II-C for the ten densest networks in the Cisco Secure Workload dataset. The absolute correlation coefficient calculated over this table between Flatness and the Fiedler Value is $0.136$; that between Flatness and Modularity is $0.568$. The low correlation with Fiedler Value ($0.136$) and the moderate correlation with Modularity ($0.568$) suggests the three measures capture distinct aspects of networks (graphs).
@@ Appendix B Bayesian Inference for High-segmentedness
To address this, we introduce a Bayesian Beta-binomial model to provide a more robust estimation of uncertainty at the boundaries [11]. The Beta distribution is the conjugate prior for the Binomial distribution, making it an ideal choice for updating our beliefs about network flatness based on sampled data.